If in a reaction HNO3 is reduced to NO2 the mass of HNO3 absorbing one mole of electrons would be -

Grade:11

1 Answers

Arun
6 years ago
We have HNO3 ------> NO, The change in oxidation number is +3; Equivalent mass of HNO3 = (1+14+ 16 +16 +16) / 3 = 63/3 = 21.0 g (around) so 21.0 g is the answer
 
Regards
Arun (askIITians forum expert)